OpenStudy (anonymous):

Sqrt(2+sqrt(2+sqrt(2+........=

OpenStudy (anonymous):

let Sqrt(2+sqrt(2+sqrt(2+........ be x

OpenStudy (anonymous):

then Sqrt(2+sqrt(2+sqrt(2+........=Sqrt(2+x)

OpenStudy (anonymous):

x=\[\sqrt{2+x}\]

OpenStudy (anonymous):

x^2 -x-2 = 0

OpenStudy (anonymous):

Not x just sqrt2+x

OpenStudy (anonymous):

I never put an x...

OpenStudy (anonymous):

wat??

OpenStudy (anonymous):

U answered question already...

OpenStudy (anonymous):

answer is 2 and -1

OpenStudy (anonymous):

Answer is sqrt(2+x), that's it...

OpenStudy (anonymous):

answer is 2 and -1 is the answer to x= sqrt(2+x)

OpenStudy (anonymous):

kkk

OpenStudy (anonymous):

I don't do numbers...

OpenStudy (anonymous):

2

OpenStudy (anonymous):

oh i am a little late i see...

OpenStudy (anonymous):

already answered

OpenStudy (zarkon):

I feel like a broken record...before you use the above 'trick' you need to establish that the process converges.

OpenStudy (anonymous):

We are all Euler here..:-)

OpenStudy (anonymous):

which in this case means you have only to give an upper bound, since it is increasing. i pick 2

OpenStudy (anonymous):

Lim sqrt(a+b(sqrt(a+b sqrt.....) = (b + sqrt(b^2 +4a))/2 (a, b>0)

OpenStudy (zarkon):

There is only one Euler ...The master of us all ;)